# Coherent synchrotron radiation by excitation of surface plasmon polariton on near-critical solid microtube surface

### v1

@@ Abstract
The resonant excitation of SPP leads to highly efficient electron acceleration and modulation as shown in Fig. 3(c) where the energy gain is $4.0\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$ in peak and the acceleration can be maintained for $40\leavevmode\nobreak\ $\mu\mathrm{m}$$ or 3 Rayleigh length $Z_{R}$. Due to the continuous trapping process, the energy spectrum is broad. Part of electrons, e.g. $18.7\leavevmode\nobreak\ $\mathrm{p}\mathrm{C}$$ or about $10\%$ of the total trapped electrons, can be accelerated to high energy $>90\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ due to the early trapping. These electrons can rotate around the channel axis, being confined closely along the interface with the helical modulation as shown in Fig. 3(a) and (b). The modulation is introduced during the injection process by the laser field with the frequency $\omega_{m}=\omega_{l}$. The rotation is driven by the azimuthal components, and the frequency $\omega_{r}$ is estimated as $\omega_{r}=\sqrt{e(E_{\phi}+\beta_{z}cB_{r})/m_{e}a}$. With the value of fields from PIC simulations, $\omega_{r}\simeq 2\pi\times 10^{13}\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$ or $\lambda_{r}=2\pi c/\omega_{r}\simeq 30\leavevmode\nobreak\ $\mu\mathrm{m}$$, as shown in Fig. 3(d). The normalised phase velocity of the modulation is given as $\beta_{ph}\simeq 1/(1-\omega_{r}/\omega_{m})>1$, which allows the constructive interference.
@@
Figure 4: (a) Radiation spectrum coherently calculated by the trajectories of the electrons from a PIC simulation in Fig. 3. Insert (b): Line plot along Valilov-Cherenkov angle, $\varphi_{vc}\simeq 0.23$, indicated by horizontal yellow dashed line. The dashed white curves are theoretical resonant frequencies of different harmonics.
@@
The helically modulated oscillation of the electron beam can result in the CSR in the form of optical shock along the Valilov-Cherenkov angle, $\cos\varphi_{vc}\simeq(1-\omega_{r}/\omega_{m})$ [11]. The radiation spectral lines are centred at frequencies $\tilde{\omega}_{n}=(n\pm 1)\omega_{r}/(1-\beta_{z}\cos\theta)$ where $n$ is an integer representing the harmonic number and $\beta_{z}$ the normalised longitudinal velocity of the beam. $\theta$ the angle with $z$-axis in $yz$-plane. For $\beta_{z}\sim 1$, $\tilde{\omega}_{n}(n=0)\simeq\omega_{m}$ along in Valilov-Cherenkov angle $\theta=\varphi_{vc}$. The radiation intensity is proportional to the square of the number of electrons inside the beam, $N$. With the parameters used in this letter, $\varphi_{vc}\simeq 0.23$. The radiation spectrum is shown in Fig. 4 (a) and (b), which is coherently calculated by using the wave optical method [24] with a virtual spherical detector in the far field where the radiation spikes appear in $\varphi_{vc}$ angle at the harmonic frequencies $\tilde{\omega}_{n}$. However, there are no advantages in modulating the bunch at the radiation frequency.
@@
Now, we theoretically consider a specific modulation, azimuthally symmetric modulation, where the initial phase of the electron trajectories is azimuthally uniform. The trajectory of an electron (denoted by $j$) inside the electron beam can be written as $\bm{r}_{j}=(x_{j},y_{j},z_{j})=a\left(\bm{\epsilon}e^{i\omega_{r}(t-t_{0j})+i\psi_{j}}+c.c.\right)/2+\bm{\eta}\beta_{z,j}c(t-t_{0j})$, where the polarization vectors are defined as $\bm{\epsilon}=(1,\pm i,0)$ and $\bm{\eta}=(0,0,1)$. $t$ is the retarded time and $t_{0j}$ the initial injection time. In general, the initial phase $\psi_{j}$ depends on the injection time as $\psi_{j}=\omega_{m}t_{0j}$, which introduces the helical modulation with azimuthal symmetry. $\omega_{m}$ is the modulation frequency and $\omega_{m}=\omega_{l}$ in our case. Now, we further consider an additional condition, $\psi_{j}=\omega_{m}t_{0j}=2\pi j/N$, which implies that the electrons inside the beam are injected uniformly in the azimuthal phase. Note that this condition does not require the electrons to be close, as is the case with conventional superradiance [12]. Since the beam is ultra-relativistic, $\beta_{z,j}\simeq\beta_{z}\simeq 1$.
[…]
$\mathcal{M}=\begin{pmatrix}-(\sin^{2}\theta\sin^{2}\phi+\cos^{2}\theta)&\sin^{2}\theta\cos\phi\sin\phi&\sin\theta\cos\theta\cos\phi\\ \sin^{2}\theta\cos\phi\sin\phi&-(\sin^{2}\theta\cos^{2}\phi+\cos^{2}\theta)&\sin\theta\cos\theta\sin\phi\\ \sin\theta\cos\theta\cos\phi&\sin\theta\cos\theta\sin\phi&-\sin^{2}\theta\end{pmatrix}\mathrm{.}$
@@
Figure 5: (a) Theoretical radiation spectrum calculated coherently (upper) and incoherently (bottom) by using an electron beam of $N=100$ electrons uniformly distributed over six modulation periods, or $4.8\leavevmode\nobreak\ $\mu\mathrm{m}$$ long. The other beam parameters are the same as those from PIC simulations, as shown in Fig. 3. Insert (b) Line plot of (a) along axis $\theta=0$ and $\phi=\pi/2$. (c) Line plot of onaxis ($\theta=0$ and $\phi=\pi/2$) spectrum in Fig. 4(a).
Using the modulated electron trajectories, the vector field of radiation can be given by
@@
with $\omega_{\pm}=N_{r}\pi(\omega_{nN}\pm\omega_{r})/\omega_{r}$, $\omega_{z}=N_{r}\pi\tilde{\omega}_{n}/\omega_{r}$ and $\omega_{nN}=\omega(1-\beta_{z,j}\sin\phi\cos\theta)+nN\omega_{r}$. $\zeta=(\tilde{\omega}a/c)(\sin\theta\sin\phi-\sin\theta\cos\phi)$. See more details in the Supplementary Material. It is seen from Eq. (3) that the radiation spectral lines are centred at frequencies $\tilde{\omega}_{n}=(nN\pm 1)\omega_{r}/(1-\beta_{z}\cos\theta\sin\phi)$, which presents the CSR in all azimuthal directions rather than being confined to the angle $\varphi_{vc}$. It also indicates that the harmonic numbers are given by $nN$. This implies the harmonics can be well isolated. For sufficiently large $N$, such as a $\mathrm{p}\mathrm{C}$-level beam, only the fundamental harmonic ($n=0$) is present in the spectrum, as shown in Fig. 5(a). This characteristic offers a significant advantage in that the precise control over the radiation spectrum can be significantly enhanced. The on-axis radiation cut-off frequency is $\tilde{\omega}_{c}\simeq 3K_{0}\gamma^{2}\omega_{r}/2$ [25]. For an average $\gamma=150$, this yields $\hbar\tilde{\omega}_{c}\simeq 31.3\leavevmode\nobreak\ $\mathrm{k}\mathrm{e}\mathrm{V}$$, which aligns with the numerical simulations presented in Fig. 5 (a) and (b).
@@
In realistic beams, such as those obtained from PIC simulations, modulated electrons are distributed over a finite radial extent of $\Delta r\simeq 0.15\leavevmode\nobreak\ $\mu\mathrm{m}$$, as shown in Fig. 3(b). This spatial spread disrupts the azimuthal symmetry, thereby impeding fully coherent radiation emission. Nonetheless, partial coherence can still be achieved. This occurs when a subset of electrons satisfies the necessary symmetry requirements, leading to an enhanced radiation intensity, as shown in Fig. 5(c), where the photon emission per electron is increased by almost two orders of magnitude compared to that predicted by incoherent calculations.
In summary, the proposed microtube can be fabricated from crystals that provide solid plasmas. A promising candidate is a nanostructured carbon nanotube forest, which offers tunable density and excellent plasma uniformity. The edge-coupling mechanism presented in this letter also enables the excitation of other modes by matching the azimuthal mode number $m$, for example, by a Laguerre-Gaussian laser [26]. The helical modulation of the electron beam can result in CSR, although it requires modulation of the fundamental radiation frequency. If the electron distribution is azimuthally uniform, the CSR can be generated in all directions. However, further investigation is required into optimising SPP excitation and acceleration to improve the azimuthal symmetry of the accelerated beam.

### v2

@@ Abstract
The resonant excitation of SPP leads to highly efficient electron acceleration as shown in Fig. 3(c), where the energy gain is $4.0\penalty 10000\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$\$ in peak and the acceleration can be maintained for $40\penalty 10000\ $\mu\mathrm{m}$\$ or 3 Rayleigh lengths $Z_{R}$. Due to the continuous trapping process, the electron beam is modulated to be helical, imprinted by the laser field. Therefore, the modulation frequency $\omega_{m}$ is $\omega_{m}=\omega_{l}$. The energy spectrum is broad. Part of electrons, e.g. $18.7\penalty 10000\ $\mathrm{p}\mathrm{C}$\$ or about $10\%$ of the total trapped electrons, can be accelerated to high energy $90\penalty 10000\ $\mathrm{M}\mathrm{e}\mathrm{V}$\$ due to the early trapping. These electrons can rotate around the channel axis while being confined closely along the interface, as shown in Fig. 3(a) and (b). The rotation is driven by the azimuthal components, and the frequency $\omega_{r}$ is estimated as $\omega_{r}=\sqrt{e(E_{\phi}+\beta_{z}cB_{r})/m_{e}a}$. With the value of fields from PIC simulations, $\omega_{r}\simeq 2\pi\times 10^{13}\penalty 10000\ $\mathrm{H}\mathrm{z}$\$ or $\lambda_{r}=2\pi c/\omega_{r}\simeq 30\penalty 10000\ $\mu\mathrm{m}$\$, as shown in Fig. 3(d). The normalised phase velocity of the modulation is given as $\beta_{ph}\simeq 1/(1-\omega_{r}/\omega_{m})>1$, which allows the constructive interference.
@@
Figure 4: (a) Radiation spectrum coherently calculated by the trajectories of the electrons from a PIC simulation in Fig. 3. Insert (b): Line plot along Vavilov-Cherenkov angle, $\varphi_{vc}\simeq 0.23$, indicated by horizontal yellow dashed line. The dashed white curves are theoretical resonant frequencies of different harmonics.
@@
The helically modulated oscillation of the electron beam can result in the CSR in the form of optical shock along the Vavilov-Cherenkov angle, $\varphi_{vc}\simeq\arccos(1-\omega_{r}/\omega_{m})$ [11]. The radiation spectral lines are centred at frequencies $\tilde{\omega}_{n}=(n\pm 1)\omega_{r}/(1-\beta_{z}\cos\theta)$ where $n$ is an integer representing the harmonic number and $\beta_{z}$ the normalised longitudinal velocity of the beam. $\theta$ is the angle with $z$-axis in $yz$-plane. For $\beta_{z}\sim 1$, $\tilde{\omega}_{n}(n=0)\simeq\omega_{m}$ along in Vavilov-Cherenkov angle $\theta=\varphi_{vc}$. The radiation intensity is proportional to the square of the number of electrons inside the beam, $N$. With the parameters used in this letter, $\varphi_{vc}\simeq 0.23$. The radiation spectrum is shown in Fig. 4 (a) and (b), which is coherently calculated by using the wave optical method [30] with a virtual spherical detector in the far field where the radiation spikes appear in $\varphi_{vc}$ angle at the harmonic frequencies $\tilde{\omega}_{n}$. The total energy radiated as CSR is $0.12\penalty 10000\ $\mathrm{m}\mathrm{J}$\$, which gives the energy conversion efficiency from the laser pulse to the CSR is $2.4\times 10^{-4}$. However, there are no advantages in modulating the bunch at the radiation frequency.
@@
Now, we theoretically consider a specific modulation, an azimuthally symmetric modulation, where the initial phase of the electron trajectories is azimuthally uniform. The trajectory of an electron (denoted by $j$) inside the electron beam can be written as $\bm{r}_{j}=(x_{j},y_{j},z_{j})=a\left(\bm{\epsilon}e^{i\omega_{r}(t-t_{0j})+i\psi_{j}}+c.c.\right)/2+\bm{\eta}\beta_{z,j}c(t-t_{0j})$, where the polarization vectors are defined as $\bm{\epsilon}=(1,\pm i,0)$ and $\bm{\eta}=(0,0,1)$. $t$ is the retarded time and $t_{0j}$ the initial injection time. In general, the initial phase $\psi_{j}$ depends on the injection time as $\psi_{j}=\omega_{m}t_{0j}$, which introduces the helical modulation with azimuthal symmetry. $\omega_{m}$ is the modulation frequency and $\omega_{m}=\omega_{l}$ in our case. Now, we further consider an additional condition, $\psi_{j}=\omega_{m}t_{0j}=2\pi j/N$, which implies that the electrons inside the beam are injected uniformly in the azimuthal phase. Note that this condition does not require the electrons to be close, as is the case with conventional superradiance [12]. Since the beam is ultra-relativistic, $\beta_{z,j}\simeq\beta_{z}\simeq 1$.
@@
Figure 5: (a) Theoretical radiation spectrum calculated coherently (upper) and incoherently (bottom) by using an electron beam of $N=100$ electrons uniformly distributed over six modulation periods, or $4.8\penalty 10000\ $\mu\mathrm{m}$\$ long. The other beam parameters are the same as those from PIC simulations, as shown in Fig. 3. Insert (b) Line plot of (a) along axis $\theta=0$ and $\phi=\pi/2$. (c) Line plot of onaxis ($\theta=0$ and $\phi=\pi/2$) spectrum in Fig. 4(a).
The radiation intensity $I$ per solid angle $d\Omega$ per frequency $d\tilde{\omega}$ in the far field can be written as $d^{2}I/d\tilde{\omega}d\Omega=\mathcal{I}_{0}\left|\bm{\mathcal{A}}\right|^{2}$, where the constant $\mathcal{I}_{0}=e^{2}\omega^{2}/4\pi^{2}c$. The radiation field $\bm{\mathcal{A}}$ is the contribution of the harmonics from the single electron and the modulation. It can be written as a superposition of the field from each single electron as $\bm{\mathcal{A}}=\sum_{j=0}^{N}\bm{\mathcal{A}}_{j}$. The radiation field from a single electron is given by $\bm{\mathcal{A}}_{j}=\int_{-\frac{N_{r}\pi}{\omega_{r}}}^{\frac{N_{r}\pi}{\omega_{r}}}\mathcal{M}\cdot\bm{\beta}_{j}(\bm{r},t)e^{i\tilde{\omega}(t-\frac{\hat{n}\cdot\bm{r}_{j}}{c})}dt$ where $\hat{n}=(\cos\phi,\sin\phi\sin\theta,\sin\phi,\cos\theta)$ is the unit vector to the observer in the coordinate system and $N_{r}$ is the number of oscillation periods, $\phi$ is the angle with the $x$-axis in the $xy$-plane. $\mathcal{M}$ is a matrix defining the spatial distribution. Using the modulated electron trajectories, the radiation field can be given by
@@
with $\omega_{\pm}=N_{r}\pi(\omega_{nN}\pm\omega_{r})/\omega_{r}$, $\omega_{z}=N_{r}\pi\tilde{\omega}_{n}/\omega_{r}$ and $\omega_{nN}=\omega(1-\beta_{z,j}\sin\phi\cos\theta)+nN\omega_{r}$. $\zeta=(\tilde{\omega}a/c)(\sin\theta\sin\phi-\sin\theta\cos\phi)$. It is seen from Eq. (2) that the radiation spectral lines are centred at frequencies $\tilde{\omega}_{n}=(nN\pm 1)\omega_{r}/(1-\beta_{z}\cos\theta\sin\phi)$, which presents the CSR in all azimuthal directions rather than being confined to the angle $\varphi_{vc}$. It also indicates that the harmonic numbers are given by $nN$. This implies the harmonics can be well separated. For sufficiently large $N$, such as a $\mathrm{p}\mathrm{C}$-level beam, only the fundamental harmonic ($n=0$) is present in the spectrum, as shown in Fig. 5(a). This is physical, as multi-wave interference depends on the number of waves. See more details in Appendix B and Supplemental Material [26]. This characteristic offers a significant advantage in that the precise control over the radiation spectrum can be significantly enhanced. The on-axis radiation critical frequency of an electron is $\tilde{\omega}_{c}\simeq 3K_{0}\gamma^{2}\omega_{r}/2$ [31]. For an average $\gamma=100$ and oscillation strength $K_{0}=10$ as seen in Fig. 3, this yields $\hbar\tilde{\omega}_{c}\simeq 6.2\penalty 10000\ $\mathrm{k}\mathrm{e}\mathrm{V}$\$, which approximately aligns with the numerical simulations presented in Fig. 5 (c).
@@
In realistic beams, such as those obtained from PIC simulations, the modulated electrons are distributed over a finite radial extent of $\Delta r\simeq 0.15\penalty 10000\ $\mu\mathrm{m}$\$, as shown in Fig. 3(b). This spatial spread disrupts the azimuthal symmetry, thereby impeding the emission of fully coherent radiation. A simple estimation of a Gaussian beam can give the form factor as $|F(\omega)|^{2}\sim e^{-\Delta r^{2}/a^{2}}$, where $\Delta r$ reduces the amplitude of $|F(\omega)|$. Nonetheless, partial coherence can still be achieved. This occurs when a subset of electrons satisfies the necessary symmetry requirement of $\psi_{j}$, leading to an enhanced radiation intensity, as shown in Fig. 5(c), where the photon emission per electron is increased by almost two orders of magnitude compared to that predicted by incoherent calculations.
In summary, in comparison to the laser plasma wakefield acceleration (LWFA)-based sources, e.g. betatron radiation, our scheme offers distinct advantages in terms of coherence, resulting in a significantly higher photon flux per electron and brightness. Although the total photon flux is below that of a free-electron laser (FEL) due to the limited charge accelerated in a single microstructure, it could offer several advantages, including micron-scale compactness, a broadband harmonic spectrum and robustness against beam parameters such as energy spread. The unique combination of these properties makes it a compelling alternative to FELs for applications where they remain impractical, such as multi-spectral coherent pump-probe experiments at the laboratory scale or attosecond-scale diagnostics requiring a broad coherent bandwidth.
The experimental feasibility of this scheme depends on several practical efforts in target fabrication, laser operation and plasma diagnostics to optimise the SPP excitation and the helical modulation of the accelerated beam. While the main challenges include actively stabilised laser-target misalignment, polarisation quality, laser contrast and jitter, and target survivability, the technologies to address these issues are readily available in the state-of-the-art laboratories. The structured microtube target can be fabricated using advanced nano techniques [32], though the surface damage will limit the high repetition rate that needs to be further investigated. Modern laser technologies can provide ultrashort optical pulses with high quality [33]. The new generation of high-resolution laboratory diagnostics can provide deep insight into physical processes of SPP excitation, electron acceleration, and radiation generation [34]. With these technique advantages, our work introduces a novel approach to generating ultra-bright synchrotron radiation.
@@ Acknowledgements.
Appendix A APPENDIX A: Dispersion relation of SPP in cylindrical surface
The dispersion relation can be derived by considering the continuity of the $E_{z}$ and $B_{\phi}$ of the SPP field at the interface of plasma and vacuum as
where $\kappa_{1}=\sqrt{k_{z}^{2}-\omega^{2}/c^{2}}$ and $\kappa_{2}=\sqrt{k_{z}^{2}-\epsilon_{2}\omega^{2}/c^{2}}$. For a large $a$, the dispersion relation in Eq. (3) resembles for $m=0$ as $k=(\omega/c)\sqrt{\epsilon_{2}/(1+\epsilon_{2})}$, which is the dispersion relation of SPP on the flat surface [16]. In the $m=0$ mode, the phase velocity $v_{ph}$ evolves similarly with that in planar geometry and only approaches $c$ as $n_{e}$ is much higher than $n_{c}$.
Appendix B APPENDIX B: Defination of matrix $\mathcal{M}$
The matrix $\mathcal{M}$ is defined as the spatial distribution of the radiation field projected to the spherical detector. It is obtained from the formula
$\hat{n}\times(\hat{n}\times\bm{\beta}_{j})e^{i\omega(t-\frac{\hat{n}\cdot\bm{r}_{j}}{c})}=\mathcal{M}\cdot\bm{\beta}_{j}e^{i\omega(t-\frac{\hat{n}\cdot\bm{r}_{j}}{c})}\mathrm{,}$ (4)
It can be written as
$\mathcal{M}$ $=\begin{pmatrix}-(\hat{n}_{y}^{2}+\hat{n}_{z}^{2})&\hat{n}_{x}\hat{n}_{y}&\hat{n}_{x}\hat{n}_{z}\\ \hat{n}_{x}\hat{n}_{y}&-(\hat{n}_{x}^{2}+\hat{n}_{z}^{2})&\hat{n}_{y}\hat{n}_{z}\\ \hat{n}_{x}\hat{n}_{z}&\hat{n}_{y}\hat{n}_{z}&-(\hat{n}_{x}^{2}+\hat{n}_{y}^{2})\end{pmatrix}$
$=\begin{pmatrix}-(\sin^{2}\theta\sin^{2}\phi+\cos^{2}\theta)&\sin^{2}\theta\cos\phi\sin\phi&\sin\theta\cos\theta\cos\phi\\ \sin^{2}\theta\cos\phi\sin\phi&-(\sin^{2}\theta\cos^{2}\phi+\cos^{2}\theta)&\sin\theta\cos\theta\sin\phi\\ \sin\theta\cos\theta\cos\phi&\sin\theta\cos\theta\sin\phi&-\sin^{2}\theta\end{pmatrix}\mathrm{.}$